OpenStudy (luigi0210):

Indefinite integral:

OpenStudy (luigi0210):

\[\LARGE \int~e^{2\theta}~sin(3\theta)~d\theta\]

zepdrix (zepdrix):

Oooo super fun problem! :)

zepdrix (zepdrix):

So this is one of those nasty integrals where you have to do Parts twice, then do a little magic to fix everything up.

zepdrix (zepdrix):

Are you at all familiar with the complex exponential?\[\Large\rm e^{i \theta}=\cos \theta+\mathcal i \sin \theta\]It makes this problem a little easier to work with. If not, that's fine. We can approach it the normal way.

OpenStudy (luigi0210):

Uh, nope I don't it >.<

OpenStudy (luigi0210):

*know it

OpenStudy (loser66):

@zepdrix you can teach us the new way, friend, Please

zepdrix (zepdrix):

Ok let's start by calling this integral something,\[\Large\rm \mathcal I= \int\limits~e^{2\theta}~\sin(3\theta)~d\theta\]This will be useful later on. With a problem like this, it doesn't matter which part you choose for u and dv. What WILL matter, is what parts you choose for your second `by parts` application.

zepdrix (zepdrix):

Since we have sine (which differentiates nicely, without giving us a negative), let's call that our u, ok?

zepdrix (zepdrix):

\[\Large\rm u=\sin(3\theta),\qquad\qquad~~\qquad dv=e^{2\theta}d \theta\]\[\Large\rm du=3\cos(3\theta)d \theta,\qquad\qquad v=\frac{1}{2}e^{2\theta}\]

zepdrix (zepdrix):

Mmmm ok that look about right so far? :o

OpenStudy (loser66):

yes, Sir :)

OpenStudy (luigi0210):

Yup, I got that too~

zepdrix (zepdrix):

\[\Large\rm \mathcal I=\frac{1}{2}e^{2\theta}\sin(3\theta)-\frac{3}{2}\int\limits e^{2\theta}\cos(3\theta)d \theta\]Mmm ok so this is what we get after our first application of IBP.

zepdrix (zepdrix):

We need to apply IBP to that integral. So the important thing here is, Since we chose the trig function for our U, we have to do that again! If we did it the other way around, we would simply undo the work we did the first time by parts.

zepdrix (zepdrix):

\[\Large\rm u=\cos(3\theta),\qquad\qquad\quad~\qquad dv=e^{2\theta}d \theta\]\[\Large\rm du=-3\sin(3\theta)d \theta,\qquad\qquad v=\frac{1}{2}e^{2\theta}\]Mmmmm ya?

OpenStudy (luigi0210):

Yes~

OpenStudy (luigi0210):

So after that we end up with a similar integral as the first, right?

zepdrix (zepdrix):

\[\rm \mathcal I=\frac{1}{2}e^{2\theta}\sin(3\theta)-\frac{3}{2}\color{royalblue}{\int\limits\limits e^{2\theta}\cos(3\theta)d \theta}\] \[\rm \mathcal I=\frac{1}{2}e^{2\theta}\sin(3\theta)-\frac{3}{2}\color{royalblue}{\left[\frac{1}{2}e^{2\theta}\cos(3\theta)+\frac{3}{2}\int\limits e^{2\theta}\sin(3\theta)d \theta\right]}\]

zepdrix (zepdrix):

Yes very good :) We end up with the integral that we started with!\[\rm \mathcal I=\frac{1}{2}e^{2\theta}\sin(3\theta)-\frac{3}{2}\color{royalblue}{\left[\frac{1}{2}e^{2\theta}\cos(3\theta)+\frac{3}{2}\mathcal I\right]}\]

zepdrix (zepdrix):

Then from there, just solve for \(\Large\rm \mathcal I\).

OpenStudy (luigi0210):

So much integration in just 3 days.. it hurts >.< But I see what you mean zepdrix, thanks again :D

zepdrix (zepdrix):

yay team \c:/

zepdrix (zepdrix):

You gotta learn the complex variable stuff!! >.< So much more fun!

sammixboo (sammixboo):

So.. Much.. Math..